OpenStudy (anonymous):

x^2 + 9x + 18, ok i got this far, no what?

OpenStudy (anonymous):

got this far from where?

OpenStudy (anonymous):

So you need to factor this polynomial

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (lgbasallote):

factor..get the answer..jump for joy

OpenStudy (anonymous):

x^2+18= -9x but no i dont know how

OpenStudy (anonymous):

x^2 + 9x + 18 =(x+3)(x+6)

OpenStudy (anonymous):

(x+6)(x+3) Notice that the 6 and 3 multiply to give positive 18, as in the orig eqn. Then 6x and 3x add up to be 9x, as in the orign eqn. Now set these two factored forms equal to 0: x+6=0 x+3=0 x=-6 x=-3 So the two solutions for this polynomial are -6, -3. Do you get that?

OpenStudy (anonymous):

remember if you got this far from x^2+18=-9x you must write x^2+9x+18=0 to create an equivalent equation x^2+9x+18 is only an expression

OpenStudy (anonymous):

Good point @Dockworker